# How Far Are We from Intelligent Visual Deductive Reasoning?

### v2

@@ 5.3 How good is the VLM’s hypothesis verification on this task?
Finally, We tested the performance of GPT-4V when it received both an oracle description and an oracle rationale. The oracle rationale, which can be found in Appendix A.1, only includes the explanation of the underlying rule without predicting the final pattern or answer. The results for 10 repetitions on manual examples are shown in Table 5 (GPT-4V + Oracle Desc. + Rationale).
Surprisingly, compared to the row representing GPT-4V + Oracle Desc., the oracle rationale did not significantly improve accuracy. In cases where the model failed, it sometimes directly generated an incorrect answer and at other times extended the rationale but still generated false answers. For example, for M-easy, GPT-4V continued to generate “the third row should have a star, as the first two boxes of the third row (square and circle) already have a star.” This indicates that hypothesis generation and verification are closely tied to deductive reasoning, and the model has not yet reached human-level performance in following hints and turning learned rules into future predictions.
@@ 5.4 How does the prompt format influence the model prediction?
We also delves into the differences in how the model performs under 0-shot and 1-shot evaluation setups. We discovered that using special sentinel tokens, such as [{BEGIN/END}_OF_EXAMPLE] , to clearly separate text prompts from images helps the model delineate task instructions from in-context examples. This method of structuring prompts is particularly effective in aiding the model’s comprehension across all tested VLMs. For instance, we show the results of GPT-4V in Table 6.
This study underscores that VLMs, unlike their text-only counterparts, can benefit significantly from a more structured format in their task prompts. Furthermore, the interaction between different modalities, such as text and image, needs to be carefully considered and evaluated.
@@ 6 Conclusion
This work is a systematic evaluation of the performance of popular Vision-Language Models (VLMs) in a variety of Raven’s Progressive Matrices (RPMs). These tasks serve as a challenging benchmark for assessing the models’ ability to reason based on visual clues. We observed that the current state-of-the-art VLMs still fall short of achieving human-level performance on these tasks, with the best-performing models being close-sourced. Our analysis of the models’ performance reveals that perceptual understanding may be the main bottleneck, as the models perform better when provided with appropriate textual descriptions. In future work, it would be intriguing to validate our hypothesis concerning the blind spot of VLMs when it comes to describing patterns. This investigation has the potential to enhance the general recognition and attentiveness capabilities of VLMs. Additionally, exploring the development of contrastive learning or reinforcement learning algorithms could further improve the model’s visual deductive reasoning abilities.
7 Acknowledgement
We thank Zijin Gu, Yusu Qian, Russ Webb, Yinfei Yang, Zhe Gan for their exceptional contribution to our work.
@@ Appendix A Appendix
A.1 Prompts for different dataset and tasks
@@
{in-context description}
{in-context rationale}
{in-context answer}
Now do the following one:
@@
{in-context description}
{in-context rationale}
{in-context answer}
Now do the following one:
@@
{in-context description}
{in-context rationale}
{in-context answer}
Now do the following one:
@@
q1: {q1 image}
q2: {q2 image}
q3: {q3 image}
q4: {q4 image}
q5: {q5 image}
q6: {q6 image}
q7: {q7 image}
q8: {q8 image}
A: {A image}
B: {B image}
C: {C image}
D: {D image}
E: {E image}
F: {F image}
For each row, analyze the changes and relationships between the images. Consider the number of shapes, the types of shapes, their positions, the shading, and any other changes that occur from one pattern to the next. Once you have identified the rule or sequence that applies to the rows, select the option (A, B, C, D, E, F) that contains the pattern which correctly completes the third row sequence.
@@
A.2 Sampling methods for model evaluations
In the main text, we present evaluation results obtained by independently running each model $10$ times and estimating their respective statistics. For GPT4-V, we found that using a zero temperature setting along with different random seeds effectively balances prediction accuracy with reasoning diversity. In the case of Gemini Pro Vision, the optimal performance is achieved at a temperature setting of $0.4$. However, for the self-consistency evaluation, we increase the sampling temperature to a maximum of $0.7$. This adjustment facilitates more varied predictions which are beneficial for majority-voting processes. For all other models, we observed that a lower temperature range, between 0 and $0.2$, is necessary. This lower temperature helps the models better adhere to instructions and improves their task prediction accuracy, though it also results in somewhat less variability in the models’ predictions.
A.3 Oracle description and rationale for each RPMs
@@
M-Medium (Oracle) Description The grid: 1. Top left box: A downward-pointing triangle with three dots in it. 2. Top center box: A leftward-pointing triangle with two dots in it. 3. Top right box: An upward-pointing triangle with one dot in it. 4. Middle left box: A rightward-pointing triangle with two dots in it. 5. Middle center box: A downward-pointing triangle with one dot in it. 6. Middle right box: A leftward-pointing triangle with three dots in it. 7. Bottom left box: An upward-pointing triangle with one dot in it. 8. Bottom center box: A rightward-pointing triangle with three dots in it. 9. Bottom right box: ’?’. The alternative shapes are: A. An upward-pointing triangle with two dots in it. B. A downward-pointing triangle with one dot in it. C. A leftward-pointing triangle with one dot in it. D. A rightward-pointing triangle with two dots in it. E. A leftward-pointing triangle with two dots in it. F. A downward-pointing triangle with two dots in it. Rationale Analyzing the grid, it appears that there’s a pattern related to the direction the triangle is pointing and the number of dots within the triangles. First, let’s establish the patterns of triangle directions and dots count: The first row has the triangles pointing downward, to the left, and then up. The second row has the triangles pointing rightward, downward, and then to the left. This implies that the direction that the triangle is pointing to is rotating clockwise in each row. Now let’s look at the pattern in the number of dots: The first row has 3, 2, 1 dots. The second row has 2, 1, 3 dots. This implies a pattern of a decreasing sequence from left to right.
@@
A.4 Rationale generated by GPT-4V
@@
In this paper, we present several examples that demonstrate the reasoning capabilities of GPT-4V for solving certain RPM tasks. From the examples, it becomes evident that as the difficulty level of the problem increases, the GPT-4V tends to struggle with its reasoning abilities and provides an explanation that is nonsensical.
Hard (rationale) Analyzing the grid, it appears that there’s a pattern along the diagonal lines from top right to bottom left direction. - The three elements in the first diagonal line (top right, middle center, bottom left) consist of a black square, a black square, and a white triangle. As we follow the diagonal line, these elements rotate 120 degrees clockwise. - The three elements in the second diagonal line (top left, middle right, bottom center) consist of a white circle, a white triangle and a black square. As we follow the diagonal line, these elements rotate 120 degrees clockwise. - This implies that the third diagonal line (top center, middle left, bottom right) should also follow this pattern.

### v3

@@ 5.3 How good is the VLM’s hypothesis verification on this task?
Finally, We tested the performance of GPT-4V when it received both an oracle description and an oracle rationale. The oracle rationale, which can be found in Appendix A.2, only includes the explanation of the underlying rule without predicting the final pattern or answer. The results for 10 repetitions on manual examples are shown in Table 5 (GPT-4V + Oracle Desc. + Rationale). Surprisingly, compared to the row representing GPT-4V + Oracle Desc., the oracle rationale did not significantly improve accuracy. In cases where the model failed, it sometimes directly generated an incorrect answer and at other times extended the rationale but still generated false answers. For example, for M-easy, GPT-4V continued to generate “the third row should have a star, as the first two boxes of the third row (square and circle) already have a star.” This indicates that hypothesis generation and verification are closely tied to deductive reasoning, and the model has not yet reached human-level performance in following hints and turning learned rules into future predictions.
@@ 5.4 How does the prompt format influence the model prediction?
We also delves into the differences in how the model performs under 0-shot and 1-shot evaluation setups. We discovered that using special sentinel tokens, such as [{BEGIN/END}_OF_EXAMPLE] to separate text prompts from images helps the model delineate task instructions from in-context examples. This method of structuring prompts is particularly effective in aiding the model’s comprehension across all tested VLMs. For instance, we show the results of GPT-4V in Table 6. Experiment results of k-shot evaluations are detailed in Appendix A.1.
This study underscores that VLMs, unlike their text-only counterparts, can benefit from a more structured format in their task prompts. Furthermore, the interaction between different modalities, such as text and image, needs to be carefully considered and evaluated.
@@ 6 Conclusion
This work is a systematic evaluation of the performance of popular Vision-Language Models (VLMs) in a variety of Raven’s Progressive Matrices (RPMs). These tasks serve as a challenging benchmark for assessing the models’ ability to reason based on visual clues. We observed that the current state-of-the-art VLMs still fall short of achieving human-level performance on these tasks, with the best-performing models being close-sourced. Our analysis of the models’ performance reveals that perceptual understanding may be the main bottleneck, as the models perform better when provided with appropriate textual descriptions. In future work, it would be intriguing to validate our hypothesis concerning the blind spot of VLMs when it comes to describing patterns. This investigation has the potential to enhance the general recognition and attentiveness capabilities of VLMs. Additionally, exploring the development of contrastive learning or reinforcement learning algorithms could further improve the model’s visual deductive reasoning abilities. To investigate whether the VLMs’ struggle is due to a lack of general reasoning or data distributional shifts, future work can be done to evaluate VLMs on pattern-based reasoning tasks with naturalistic images, and investigate the impact of additional instruction fine-tuning on RPMs.
@@ Appendix A Appendix
A.1 K-shot evaluation
We conduct the K-shot evaluations with GPT-4V, where K ranges from 0 to 4. The demonstrations are annotated by humans. The testing data is a subset of the Raven dataset, from which we select examples of the simplest mode (20 examples). For each K, we run the evaluation 4 times. The results are shown in Figure 4.
Figure 4: K-shot evaluation results with Raven subset.
From this figure, it is evident that few-shot evaluation generally outperforms zero-shot evaluation. However, when K is small, there is a large variance observed. As K increases, the mean accuracy slightly improves, and the variance significantly decreases. This trend is similar to observations in text-based large language models Zhao et al. (2021).
A.2 Prompts for different dataset and tasks
@@
In Section 5.4, we explain why we chose the particular order of the prompt and query image order. We also provide the other prompts we used in below.
@@
{in-context description} {in-context rationale} {in-context answer} Now do the following one:
@@
{in-context description} {in-context rationale} {in-context answer} Now do the following one:
@@
{in-context description} {in-context rationale} {in-context answer} Now do the following one:
@@
q1: {q1 image} q2: {q2 image} q3: {q3 image} q4: {q4 image} q5: {q5 image} q6: {q6 image} q7: {q7 image} q8: {q8 image} A: {A image} B: {B image} C: {C image} D: {D image} E: {E image} F: {F image} For each row, analyze the changes and relationships between the images. Consider the number of shapes, the types of shapes, their positions, the shading, and any other changes that occur from one pattern to the next. Once you have identified the rule or sequence that applies to the rows, select the option (A, B, C, D, E, F) that contains the pattern which correctly completes the third row sequence.
@@
A.3 Sampling methods for model evaluations
In the main text, we present evaluation results obtained by independently running each model $10$ times and estimating their respective statistics. For GPT4-V, we found that using a zero temperature setting along with different random seeds effectively balances prediction accuracy with reasoning diversity. In the case of Gemini Pro Vision, the optimal performance is achieved at a temperature setting of $0.4$. However, for the self-consistency evaluation, we increase the sampling temperature to a maximum of $0.7$. This adjustment facilitates more varied predictions which are beneficial for majority-voting processes. For all other models, we observed that a lower temperature range, between 00 and $0.2$, is necessary. This lower temperature helps the models better adhere to instructions and improves their task prediction accuracy, though it also results in somewhat less variability in the models’ predictions.
A.4 Oracle description and rationale for each RPMs
@@
M-Medium (Oracle) Description The grid: 1. Top left box: A downward-pointing triangle with three dots in it. 2. Top center box: A leftward-pointing triangle with two dots in it. 3. Top right box: An upward-pointing triangle with one dot in it. 4. Middle left box: A rightward-pointing triangle with two dots in it. 5. Middle center box: A downward-pointing triangle with one dot in it. 6. Middle right box: A leftward-pointing triangle with three dots in it. 7. Bottom left box: An upward-pointing triangle with one dot in it. 8. Bottom center box: A rightward-pointing triangle with three dots in it. 9. Bottom right box: ’?’. The alternative shapes are: A. An upward-pointing triangle with two dots in it. B. A downward-pointing triangle with one dot in it. C. A leftward-pointing triangle with one dot in it. D. A rightward-pointing triangle with two dots in it. E. A leftward-pointing triangle with two dots in it. F. A downward-pointing triangle with two dots in it. Rationale Analyzing the grid, it appears that there’s a pattern related to the direction the triangle is pointing and the number of dots within the triangles. First, let’s establish the patterns of triangle directions and dots count: The first row has the triangles pointing downward, to the left, and then up. The second row has the triangles pointing rightward, downward, and then to the left. This implies that the direction that the triangle is pointing to is rotating clockwise in each row. Now let’s look at the pattern in the number of dots: The first row has 3, 2, 1 dots. The second row has 2, 1, 3 dots. This implies a pattern of a rotation of a decreasing sequence.
@@
A.5 Description generated by GPT-4V
We provide several additional sampled descriptions in the table below that illustrate the perceptual blind spot in describing RPM problems. GPT-4V encounters challenges in recognizing and differentiating intricate spatial relationships and patterns in the M-Medium and M-Hard problem sets.
M-Medium (description - additional samples #2) The grid: 1. Top left box: Inverted triangle with three dots clustered towards the bottom. 2. Top center box: Upright triangle with three dots clustered towards the narrow point. 3. Top right box: Inverted triangle with one dot in the center. 4. Middle left box: Upright triangle with three dots forming a line along the base. 5. Middle center box: Inverted triangle with two dots, aligned vertically in the center. 6. Middle right box: Upright triangle with two dots, one at the tip and one at the center of the base. 7. Bottom left box: Inverted triangle with two dots forming a line along the base. 8. Bottom center box: ”?”. The alternative shapes are: A. Upright triangle with two dots side by side at the base. B. Inverted triangle with one dot at the center of the base. C. Upright triangle with one dot in the center. D. Inverted triangle with two dots side by side near the tip. E. Upright triangle with two dots, one at the center of the base and one at the tip. F. Upright triangle with three dots, forming a line along the base.
M-Medium (description - additional samples #3) The grid: 1. Top left box: An inverted triangle with three black dots forming a vertical line at the center. 2. Top center box: An upright triangle with three black dots forming a vertical line at the center. 3. Top right box: An inverted triangle with a single black dot at the center. 4. Middle left box: An upright triangle with two black dots forming a horizontal line at the center. 5. Middle center box: An inverted triangle with two black dots forming a horizontal line at the center. 6. Middle right box: An upright triangle with a single black dot at the center. 7. Bottom left box: An inverted triangle with one black dot at a bottom vertex. 8. Bottom center box: ”?”. The alternative shapes are: A. An upright triangle with two black dots forming a vertical line at the center. B. An upright triangle with a single black dot at the center. C. An inverted triangle with a single black dot at a bottom vertex. D. An inverted triangle with two black dots forming a horizontal line at the center. E. An upright triangle with two black dots forming a horizontal line at the center. F. An inverted triangle with three black dots forming a horizontal line at the center.
M-Hard (description - additional samples #1) The grid: 1. Top left box: hollow circle, hollow triangle, solid square. 2. Top center box: solid circle, hollow square, hollow triangle. 3. Top right box: solid square, solid triangle, hollow circle. 4. Middle left box: solid circle, hollow triangle, hollow square. 5. Middle center box: solid square, solid triangle, solid circle. 6. Middle right box: hollow square, hollow circle, solid triangle. 7. Bottom left box: hollow triangle, solid square, solid circle. 8. Bottom center box: solid circle, hollow triangle, hollow square. 9. Bottom right box: ”?”. The alternative shapes are: A. Hollow circle, hollow triangle, solid square. B. Solid circle, hollow square, hollow triangle. C. Hollow circle, hollow square, solid triangle. D. Hollow circle, hollow triangle, solid square. E. Solid square, hollow triangle, hollow circle. F. Hollow square, hollow triangle, solid circle.
M-Hard (description - additional samples #2) The grid: 1. Top left box: circle at the top, triangle in the middle, square at the bottom. 2. Top center box: triangle at the top, circle in the middle, small square at the bottom. 3. Top right box: square at the top, triangle in the middle, circle at the bottom. 4. Middle left box: circle at the top, square in the middle, triangle at the bottom. 5. Middle center box: square at the top, two triangles in the middle, circle at the bottom. 6. Middle right box: triangle at the top, circle in the middle, square at the bottom. 7. Bottom left box: triangle at the top, two squares in the middle, circle at the bottom. 8. Bottom center box: square at the top, circle in the middle, two triangles at the bottom. 9. Bottom right box: ”?”. The alternative shapes are: A. Circle at the top, triangle in the middle, square at the bottom. B. Circle at the top, square in the middle, two triangles at the bottom. C. Circle at the top, small square in the middle, triangle at the bottom. D. Two circles at the top, square in the middle, triangle at the bottom. E. Square at the top, circle in the middle, triangle at the bottom. F. Triangle at the top, square in the middle, circle at the bottom.
A.6 Rationale generated by GPT-4V
@@
We present several examples that demonstrate the reasoning capabilities of GPT-4V for solving certain RPM tasks. From the examples, it becomes evident that as the difficulty level of the problem increases, the GPT-4V tends to struggle with its reasoning abilities and provides an explanation that is nonsensical.